OpenStudy (anonymous):

The inequality 2x+10

OpenStudy (zarkon):

1/2

OpenStudy (anonymous):

what is b and how did you get that?

OpenStudy (zarkon):

\[x+11 \leq3x+10\] \[1\leq 2x\]...

OpenStudy (anonymous):

what about the 2x+10

OpenStudy (anonymous):

b=1?

OpenStudy (zarkon):

\[2x+10 \leq 3x+10\] \[2x\leq3x\] \[0\leq x\] \[0\leq x\text{ and }1/2\leq x\] gives \[1/2\leq x\]

OpenStudy (zarkon):

b=1

OpenStudy (anonymous):

Shouldnt you use the x+11?

OpenStudy (zarkon):

yes...I did

OpenStudy (anonymous):

how do i know which one is A and which one is B

OpenStudy (zarkon):

A<B